A father has two sons who were born three years apart. In 20 years from now, the sum of the ages of the sons will be equal to their father's then. Find the present age of the younger son if, when he was born, the father was nine times as old as his elder son
Answer by ankor@dixie-net.com(22740)

A father has two sons who were born three years apart.
In 20 years from now, the sum of the ages of the sons will be equal to their father's then.
Find the present age of the younger son if, when he was born, the father was nine times as old as his elder son.
:
let the two sons be s and (s+3)
let f = the fathers age
"In 20 years from now, the sum of the ages of the sons will be equal to their father's then."
f + 20 = (s+20) + (s+3+20)
f + 20 = 2s + 43
f = 2s + 43 - 20
f = 2s + 23
" Find the present age of the younger son if, when he was born, the father was nine times as old as his elder son"
We know when the younger son was born, the older son was 3 yrs old, therefore:
9 * 3 = 27 yrs was the father's age then
Now we can say fathers age now
f = s + 27
In the first equation replace f with s + 27
s + 27 = 2s + 23
27 - 23 = 2s - s
4 = s
The younger son is 4 yrs old, the older son is 7, father is 31
:
:
See if that checks in the statement:
" In 20 years from now, the sum of the ages of the sons will be equal to their father's then. "
31 + 20 = (4+20) + (7+20)
51 = 24 + 27